A pharmaceutical company claims that its new drug reduces systolic blood pressure .

Using the date find the 90% confidence interval for the true difference in blood pressure for each patient taking the new drug. Assume that the blood pressures are normally distributed for the population of patients both before and after taking the new drug.

Blood pressure before 199, 181, 153, 168, 202, 194, 175, 191, 189

Blood pressure after 173, 162, 144, 156, 193, 170, 166, 176, 171

Construct the 90% confidence interval.
Answer #1
Before After Difference
199 173 26
181 162 19
153 144 9
168 156 12
202 193 9
194 170 24
175 166 9
191 176 15
189 171 18

∑d = 141

∑d² = 2549

n = 9

Mean , x̅d = Ʃd/n = 141/9 = 15.6667

Standard deviation, sd = √[(Ʃd² - (Ʃd)²/n)/(n-1)] = √[(2549-(141)²/9)/(9-1)] = 6.5192

90% Confidence interval :

At α = 0.1 and df = n-1 = 8, two tailed critical value, t-crit = T.INV.2T(0.1, 8) = 1.860

Lower Bound = x̅d - t-crit*sd/√n = 15.6667 - 1.86 * 6.5192/√9 = 11.626

Upper Bound = x̅d + t-crit*sd/√n = 15.6667 + 1.86 * 6.5192/√9 = 19.708

11.626 < µd < 19.708